OpenStudy (anonymous):

The total cost for 8 bracelets, including shipping was $54. The shipping charge was $6. Determine the cost for each bracelet. Show your work

OpenStudy (anonymous):

well the equation you have was 8x + 6 = 54 So first you would subtract 6 from both sides. - 6 -6 ------------ 8x = 48 Now divide 8 from both sides. /8 /8 ------- x = 6

OpenStudy (anonymous):

thanks